OpenStudy (anonymous):

For 0°<θ<90°, the maximum value of 2/[3+sin^(2)θ] is ___?

OpenStudy (anonymous):

\[\frac{ 2 }{ 3+\sin ^{2} \theta }\]

Directrix (directrix):

What value of theta makes this denominator have its smallest possible value over the interval [0,90] ? In short, at what value of theta is sin^(2)θ a minimum over the given interval? Denominator: [3+sin^(2)θ]

OpenStudy (anonymous):

for the given function to be maximum sin theta shuld be minimum

hartnn (hartnn):

anyone can help :)

OpenStudy (anonymous):

30°?

hartnn (hartnn):

in the interval, 0°<θ<90°, you say sin is minimum at 30 ?? please reconsider....

OpenStudy (anonymous):

isn't the max?

OpenStudy (anonymous):

wait, i don't understand the denominator

hartnn (hartnn):

to get maximum value of 2/[3+sin^(2)θ] , you'll need minimum of denominator [3+sin^(2)θ] , right ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

60?

hartnn (hartnn):

to get minimum of denominator [3+sin^(2)θ] you need minimum of sin^2 theta, right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so what is minimum of sin in the interval 0°<θ<90°, ?

OpenStudy (anonymous):

1......?

OpenStudy (anonymous):

umm, at start i have the answer 1/2 but i don't know is it correct

hartnn (hartnn):

what about 0 degrees, what is sin 0 ? is that minimum ?

OpenStudy (anonymous):

0°<θ<90° not 0°≤θ<90°

OpenStudy (anonymous):

sin0° = 0

hartnn (hartnn):

ok, but theta can be very very near to 0, denoted by \(\theta \rightarrow 0^+\) which means, theta \(\ne0\) but very near to 0 and greater than 0. so, you can use the value of theta = 0 to find the minimum

hartnn (hartnn):

then just put sin theta =0 in your expression...

OpenStudy (anonymous):

2/3?

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

thank you @hartnn :)